OpenStudy (miss.rose):

Can someone help me with some questions? Please I will medal!!!!!! And Fan!!!!!!!!

OpenStudy (anonymous):

i will

OpenStudy (shiburinga):

I wll try :)

OpenStudy (miss.rose):

Okay. Here is the first question: Determine the result of 10 + –4. How far away on the number line will this result be from 10 and in what direction? P.S there are eight questions.

OpenStudy (shiburinga):

Lol sorry am back :)

OpenStudy (shiburinga):

What do you think it is..?

OpenStudy (miss.rose):

The numbers are different, one is a positive while the other one is a negative. So this problem can be changed to a subtraction problem. So start at 10 and subtract 4 to land on 6. Also, because 10 is a positive number we would start on the right side of the number line and move 4 places left and we would land on 6. With that being said 10 + (-4) = 6. Okay. Just so you know this is where I right my own answer. And I am open to suggestions. ^_^

OpenStudy (shiburinga):

Okay xP

OpenStudy (shiburinga):

agree with ur answer xP

OpenStudy (miss.rose):

Ok. Next question: Create a real-world problem that represents 2/3 and 2/5 . My answer: George ran 2/3 at the jog-a-thong. But his friend ran 2/5 at the jog-a-thong. How many laps did they complete all together? First, we need a common denominator. Which is 15 because both numbers go in to it evenly. Next, we need to correct the numerators because we changed the denominator. So 2/3 will become 10/15 and 2/5 will become 6/15. Then, we add these two fractions together to get 16/15 which can be turned in to a mixed number of 1 1/15. Lastly, our finally answer is 1 1/15.

OpenStudy (shiburinga):

Mmm you might want to change "jog-a-thong" to "jog-a-thon" xD Not sure if you know what a thong is, but.. you might want to change it. xD

OpenStudy (shiburinga):

Other than that I agree :)

OpenStudy (miss.rose):

Ok. Thanks! Next Question: Create a problem that uses the associative property to find the sum of at least 3 rational numbers. Show how the associative property is used to solve the problem. My answer: Associative property switches the order of groupings. A problem that uses the associative property is: 23 + (7 + 5)=(23 + 7) + 5 So first we do the grouping which is (7 + 5) this equals 12. Next, we want to add 23 and 12 together which give us the answer of 35. So the answer to the first part is 35. Next Part: First, we do the grouping of (23 + 7) which equals 30. Next, we add 30 and 5 which gives us the answer of 35. So both answers stay the same just different grouping.

OpenStudy (shiburinga):

Lol I agree with that to :o

OpenStudy (miss.rose):

Okay. Next Question: Apply the definition of subtraction and properties of operations to find the result. 1/9 - 4/5 - 1/9 My answer is: I need a little help on that one.

OpenStudy (shiburinga):

Well.. we first need to find a Common Denominator.

OpenStudy (miss.rose):

So 45?

OpenStudy (shiburinga):

Yep xP

OpenStudy (shiburinga):

So now we multiple each denominator by the number it needs to get to 45, thne do the same the top. :)

OpenStudy (miss.rose):

1/9 = 5/45 4/5 = 36/45

OpenStudy (shiburinga):

Mhm.

OpenStudy (shiburinga):

Now we need to subtract. :)

OpenStudy (miss.rose):

OMG! I am going to die!!!! My screen just froze while I was in the middle of the test! When it reloaded it logged me out of the test!

OpenStudy (shiburinga):

:O

OpenStudy (shiburinga):

THAT IS SO WORNG

OpenStudy (shiburinga):

WRONG*

OpenStudy (miss.rose):

And I was only half way done!!! I am going to go ask my Mom to email the teacher. BRB

OpenStudy (shiburinga):

Ok :3

OpenStudy (miss.rose):

Okay. I am back. But that was the last question and I pretty much figured it. THANK YOU FOR ALL YOUR HELP!!!! And I mean all!!!

OpenStudy (shiburinga):

Thanks and yw :)